# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v5

@@ 3 Motivation
In current DRAM chips, the Memory Controller (MC) is in charge of managing DRAM maintenance operations such as periodic refresh, RowHammer protection, and memory scrubbing. When DRAM vendors make modifications in a DRAM maintenance mechanism, the changes often need to be reflected to the MC design as well as the DRAM interface, which makes such modifications very difficult. As a result, implementing new or modifying existing maintenance operations requires multi-year effort by multiple parties that are part of the JEDEC committee. A prime example to support our argument is the most recent DDR5 standard [53], which took almost a decade to develop after the initial release of DDR4. DDR5 introduces changes to key issues we study in this paper: DRAM refresh, RowHammer protection, and memory scrubbing. We discuss these changes as motivating examples to show the shortcomings of the status quo in DRAM.
@@ 3.1 DRAM Refresh
DDR5 introduces Same Bank Refresh (SBR), which refreshes one bank in each bank group at a time instead of simultaneously refreshing all banks as in DDR4 [50, 96]. SBR improves bank availability as the MC can access the non-refreshing banks while certain banks are being refreshed. DDR5 implements SBR by introducing a new REFsb command [53]. This new command necessitates changes in the DRAM interface and in the MC design.
@@ 3.2 RowHammer Protection
In DDR4, DRAM vendors implement in-DRAM RowHammer protection mechanisms by performing Targeted Row Refresh (TRR) operations within the slack time available when performing regular refresh. However, prior works have shown that in-DRAM TRR is vulnerable to certain memory access patterns [41, 31, 48]. DDR5 specifies a Refresh Management (RFM) mechanism that an MC implements to aid the RowHammer protection implemented in DRAM chips. As part of RFM, the MC uses a set of counters to keep track of row activations to each DRAM bank. When a counter reaches a specified threshold value, the MC issues the new RFM command to DRAM chips. A DRAM chip then internally performs an undocumented operation with the time allowed for RFM to mitigate the RowHammer effect. Since the first RowHammer work [70] in 2014, it took about 7 years to introduce DDR5 RFM, which requires significant changes in the DRAM interface and the MC design. Still, RFM is likely not a definitive solution for RowHammer as it does not outline a specific RowHammer defense with security proof. Rather, RFM provides additional time to a DRAM chip for internally mitigating the RowHammer effect.
@@ 4 Self-Managing DRAM
We introduce the Self-Managing DRAM (SMD) architecture.
@@ 4.1 Overview of SMD
SMD has a flexible interface that enables efficient implementation of multiple DRAM maintenance operations within the DRAM chip. Our key insight is to prevent the MC from accessing a DRAM row that is under maintenance by prohibiting the MC from activating the row until the maintenance is complete. The key idea of SMD is to provide a DRAM chip with the ability to reject an ACT command via a single-bit negative-acknowledgment (ACT_NACK) signal. An ACT_NACK informs the MC that the DRAM row it tried to activate is under maintenance, and thus temporarily unavailable. Leveraging the ability to reject an ACT, a maintenance operation can be implemented completely within a DRAM chip.
As Fig. 2 shows, SMD preserves the general DRAM interface222We consider the widely-available DDR4 as a baseline DRAM and explain the changes required to implement SMD on top of DDR4. However, other DRAM interfaces (e.g., DDR5, LPDDR5, GDDR6, HBM2) can also be modified in a similar way to support SMD. and only adds a single uni-directional pin to the physical interface of a DRAM chip. The extra pin is used for transmitting the ACT_NACK signal from the DRAM chip to the MC. Upon receiving an ACT_NACK, the MC notices that the previous row activation is rejected, and thus it need to be re-issued at later time. While introducing changes in the MC to handle ACT_NACK, SMD also simplifies MC design and operation as the MC no longer implements control logic to periodically issue DRAM maintenance commands. For example, the MC does not 1) prepare a bank for refresh by precharging it, 2) implement timing parameters relevant to refresh, nor 3) issue REF commands. The MC still maintains the bank state information (e.g., whether and which row is open in a bank) and respects the DRAM timing parameters associated with commands for performing access (e.g., ACT, PRE, RD, WR).
@@
Fig. 3 shows the organization of a bank in an SMD chip. SMD divides the rows in a DRAM bank into multiple lock regions. SMD provides the number and the size of the lock regions to the MC via DRAM mode registers [50]. For each lock region, SMD stores an entry in a Lock Region Table (LRT) to indicate whether or not a lock region is reserved for performing a maintenance operation.
A maintenance operation and an ACT can be performed concurrently in the same bank on different lock regions. To enable this, SMD implements a row address latch in order to drive two local row decoders with two independent row addresses.333Having a row address latch per lock region enables two or more maintenance operations to concurrently happen on different lock regions within a bank. However, to keep our design simple, we restrict a maintenance operation to happen on one lock region while the MC can access another lock region. Our design can be easily extended to increase concurrency of maintenance operations across multiple lock regions. When a maintenance operation and an ACT arrive at the same lock region at the same time, the multiplexer shown in Fig. 3 prioritizes the maintenance operation and the SMD chip rejects the ACT.
@@
For a maintenance operation to take place, the Lock Controller first sets the lock bit in the LRT entry that corresponds to the lock region on which the maintenance operation is to be performed. When the MC attempts to open a row in a locked region, the Lock Controller generates an ACT_NACK.
4.2 Handling an ACT_NACK Signal
@@
Fig. 4 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It keeps re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. While waiting for ARI, the MC can attempt activating a row from a different lock region or bank, to overlap the ARI latency with a useful operation.
4.2.1 Setting the ACT Retry Interval (ARI)
@@
Setting ARI to a very low or high value relative to the expected duration of the maintenance operations can have a negative impact on system performance and energy efficiency. When too low, the MC frequently re-issues an ACT, and the DRAM chip rejects many of these ACT commands. The rejected ACT commands waste energy and potentially delay other row activations that the MC could have issued to serve other memory requests. When ARI is too high, the MC may delay re-issuing a rejected ACT for too long, and the ACT gets delayed significantly even after the completion of the maintenance. We empirically find that is a favorable configuration for the five maintenance mechanisms that we evaluate in this work (section 5).
4.2.2 Latency
@@
An SMD chip sends an ACT_NACK (in case an ACT command is to be rejected) DRAM command bus cycles after receiving the ACT. Thus, the MC knows that an ACT is successful when it does not receive an ACT_NACK in cycles after issuing the ACT. It is desirable for to be as low as possible so that the MC can be quickly notified when an ACT fails, and it can attempt to activate a different row in the same bank while waiting for ARI to retry the rejected ACT. Further, if is larger than tRCD, the latency of RD and WR commands increases (by ) as the MC should not issue a RD or WR without ensuring that the row is successfully opened.
@@
We estimate the overall latency based on the latency of an RD in a conventional DRAM. An RD has a latency breakdown that resembles the latency breakdown of a . An RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [50], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Table (LRT) is (section 7.6). Given the relatively low complexity of the LRT compared to the datapath that is involved during an RD, we believe the second component of the latency breakdown and thus the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations, we find that small latencies (e.g., ) have negligible effect on system performance mainly because the number of rejected ACTs constitute a small portion of all ACTs.
4.2.3 ACT_NACK Divergence Across DRAM Chips
@@
SMD maintenance operations take place independently in each DRAM chip. Therefore, when a DRAM rank, which can be composed of multiple DRAM chips operating in lock step444DDRx systems typically consist of DRAM ranks built with multiple chips (e.g., eight chips with 8-bit data bus in each chip), whereas LPDDRx systems typically consist of DRAM ranks built with a single DRAM chip., receives an ACT command, some of the chips may send an ACT_NACK while others do not. As a result of this divergence, the row becomes partially open in chips that do not send ACT_NACK.
As a solution, the MC can take one of the following two approaches. First, it can issue a PRE command to close the partially open row. Upon receiving the PRE, chips that sent an ACT_NACK and do not have an open row can simply ignore the PRE. The advantage of this approach is that, after the PRE, the MC can attempt to open another row from the same bank. However, as a downside, precharging the partially activated row would waste energy. As a second approach, the MC can wait for ARI and re-issue the same ACT command to attempt activating the row in the chips that previously sent an ACT_NACK signal. This approach does not waste row activation energy but it prevents the MC from attempting to activate another row in the same bank while waiting for ARI. We empirically find that the second approach performs slightly better (by 0.04% on average) mainly because the benefits of the first approach are limited as a partially activated row cannot be immediately precharged due to tRAS, which constitutes a large part of ARI555By default, we set while tRAS is about [96].).
Ensuring forward progress for memory requests. SMD does not cause deadlocks, livelocks or other starvation issues because of the following three characteristics. We use a memory controller design that periodically reissues the ACT command to an ACT_NACK’d region until the command does not get ACT_NACK’d to serve the highest priority memory request. If a high-priority request is ACT_NACK’d, it is in the memory request scheduling algorithm’s best interest to keep issuing ACT commands to the locked region until the lock is released (as opposed to scheduling another request with a lower priority to a potentially not-locked region). Based on this memory controller design, we explain how SMD prevents resource starvation issues. First, a DRAM chip remains idle for ARI between the end of a maintenance operation and the start of the next maintenance operation (Section 4.2). Second, the MC periodically re-issues an ACT every ARI to serve the memory request with the highest priority. Third, a DRAM chip cannot start performing a maintenance operation when one of its rows is active. If a DRAM chip issues an ACT_NACK command, the periodically issued (every ARI) ACT commands will activate the row in this chip after the maintenance operation ends. Meanwhile, other DRAM chips that did not issue an ACT_NACK command cannot start issuing ACT_NACK commands because these chips already have the DRAM row activated. Thus, the maximum memory request stall time is the time it takes a DRAM chip to perform the most time taking maintenance operation. To limit the maximum potential memory request stall time, 1) long time taking maintenance operations can be divided into many smaller short time taking maintenance operations, 2) the maximum time a maintenance operation takes can be specified in the DRAM standard. In our design, the most time taking maintenance operation is SMD-FR (Section 5.1.1), which takes to refresh 16 DRAM rows in a lock region.
4.3 Region Locking Mechanism
SMD locks rows at a lock region granularity.
4.3.1 Locking a Region
A maintenance operation can be performed on a lock region only after locking it. Therefore, a maintenance mechanism must lock the region that includes the rows that should undergo a maintenance. A maintenance mechanism can only lock a region that is not already locked. This is to prevent different maintenance mechanisms from interfering with each other. Two maintenance operations can concurrently proceed if they target different lock regions.666An ECC scrubbing operation (Section 5.3) uses the data bus, which is a shared resource across all regions. Therefore, a scrubbing operation locks all regions at once to prevent any other maintenance operation or memory access from happening at the same time. Once having locked a region, the region remains locked until the completion of the maintenance. Afterward, the maintenance mechanism releases the lock to allow 1) the MC to access the region and 2) other maintenance mechanisms to lock the same region.
4.3.2 Lock Region Size
A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size such that a lock region spans one or multiple subarrays. This is because a maintenance operation uses the local row buffers in a subarray, and thus having a subarray shared by multiple lock regions will cause a conflict when accessing a row in a different lock region that maps to the subarray under maintenance. We design and evaluate SMD assuming a default lock region size of 512-row subarrays. However, a lock region can be designed to span fewer than subarrays or it can span the entire bank.
Modern DRAM chips typically use the density-optimized open-bitline architecture [1, 47], which places sense amplifiers both on top and bottom sides of a subarray and adjacent subarrays share sense amplifiers. With the open-bitline architecture, the MC should not be allowed to access a row in a subarray that is adjacent to one under maintenance. To achieve this, the Lock Controller simply sends an ACT_NACK when the MC attempts to activate a row in a subarray adjacent to a subarray in a locked region. Consequently, when SMD locks a region that spans 512-row subarrays, it prevents the MC from accessing subarrays in a bank with rows ( subarrays). In the folded-bitline architecture [1, 47], adjacent subarrays do not share sense amplifiers with each other. Therefore, SMD prevents access only to the subarrays in a locked region. We evaluate SMD using the density-optimized open-bitline architecture.
4.3.3 Ensuring Timely Maintenance Operations
A maintenance mechanism cannot lock a region that has an active DRAM row. To lock the region, the maintenance mechanism must wait for the MC to precharge the row. The MC may keep a DRAM row active for too long, which may delay a maintenance operation to a point that affects DRAM reliability (e.g., delaying a refresh operation too long as such a retention failure occurs). To prevent this, we rely on the existing time limit for a row to remain open, i.e., tRAS has an upper limit of 9x tREFI [50, 96, 45]).

### v6

@@ 3 Motivation
In current DRAM chips, the MC is in charge of managing DRAM maintenance operations such as periodic refresh, RH protection, and memory scrubbing. When DRAM vendors modify a DRAM maintenance mechanism, the changes often need to be reflected to the MC design as well as the DRAM interface, which makes such modifications very difficult. As a result, implementing new or modifying existing maintenance operations, no matter how fast such implementations or modifications could be developed, can only be realized after a multi-year effort by multiple parties that are part of the JEDEC committee. A prime example to support our argument is the most recent DDR5 standard [72], which took almost a decade to develop after the initial release of DDR4. Even though it might not have taken DRAM designers a decade to develop new maintenance mechanisms, the mechanisms are implemented only after the new standard is released. DDR5 also introduces changes to key issues we study in this paper: DRAM refresh, RH protection, and memory scrubbing. We discuss these changes as motivating examples to show the shortcomings of the status quo in DRAM.
@@ 3.1 DRAM Refresh
DDR5 introduces Same Bank Refresh (SBR), which refreshes one bank in each bank group at a time instead of refreshing all banks as in DDR4 [78, 83]. SBR improves bank availability as the MC can access the non-refreshing banks while certain banks are being refreshed. DDR5 implements SBR with a new REFsb command [72]. Implementing REFsb necessitates changes in the DRAM interface and MC.
@@ 3.2 RowHammer Protection
In DDR4, DRAM vendors implement in-DRAM RowHammer protection mechanisms by performing Targeted Row Refresh (TRR) within the slack time available when performing regular refresh. However, prior works have shown that in-DRAM TRR is vulnerable to certain memory access patterns [74, 75, 76]. DDR5 specifies the Refresh Management (RFM) that an MC implements to aid in-DRAM RowHammer protection. As part of RFM, the MC uses counters to keep track of row activations to each DRAM bank. When a counter reaches a specified threshold value, the MC issues the new RFM command to DRAM chips. A DRAM chip then performs an undocumented operation with the time allowed for RFM to mitigate RowHammer. Since the first RowHammer work [10] in 2014, it took about 7 years to introduce DDR5 RFM, which requires significant changes in the DRAM interface and the MC design. Still, RFM is likely not a definitive solution for RowHammer as it does not outline a RowHammer defense with security proof. Rather, RFM provides additional time to a DRAM chip for internally mitigating the RowHammer.
@@ 4 Self-Managing DRAM
To enable autonomously and efficiently performing maintenance operations in a DRAM chip, we introduce the Self-Managing DRAM (SMD) architecture. SMD is part of a continuum between fully master-slave and fully request-reply interfaces. It tries to achieve a balance that enables faster innovation and easier adoption of new ideas.
@@ 4.1 Overview of SMD
SMD has a flexible interface that enables efficient implementation of multiple DRAM maintenance operations within the DRAM chip. The key idea of SMD is to provide a DRAM chip with the ability to reject an ACT command via a single-bit negative-acknowledgment (ACT_NACK) signal. An ACT_NACK informs the MC that the DRAM row it tried to activate is under maintenance, and thus temporarily unavailable. Leveraging the ability to reject an ACT, a maintenance operation can be implemented completely within a DRAM chip.
Organization. SMD preserves the general DRAM interface and uses a single uni-directional pin (which already exists in DDR4/5, see section 7.4) in the physical interface of a DRAM chip. This pin is used for transmitting the ACT_NACK signal from the DRAM chip to the MC.
@@
Fig. 2 shows the organization of a bank in an SMD chip. SMD divides a DRAM bank into multiple lock regions of equal size. SMD stores an entry in a per-bank Lock Region Table (LRT) to indicate whether or not a lock region is reserved for performing a maintenance operation.
@@
Operation. A maintenance operation takes place as the Lock Controller sets the lock bit in the LRT entry that corresponds to the lock region on which the maintenance operation is to be performed. When the MC attempts to open a row in a locked region, the Lock Controller generates an ACT_NACK.
A maintenance operation and an ACT can be performed concurrently in the same bank on different lock regions.333An ECC scrubbing operation (section 5.3) uses the data bus, which is a shared resource across all regions. Therefore, a scrubbing operation reservers all lock regions inside a bank at once to prevent any other maintenance operation or memory access from happening at the same time. To enable this, SMD implements a row address latch in order to drive two local row decoders with two independent row addresses.444One row address latch per lock region enables two or more maintenance operations to concurrently happen on different lock regions within a bank. To keep our design simple, we restrict a maintenance operation to happen in one lock region while the MC accesses another lock region. Our design can be easily extended to increase concurrency of maintenance operations across multiple lock regions. When a maintenance operation and an ACT arrive at the same lock region at the same time, the MUX prioritizes the maintenance operation and the SMD chip rejects the ACT.
4.2 Region Locking Mechanism
A maintenance operation can be performed on a lock region only after locking it. Therefore, a maintenance mechanism must lock the region that includes the rows that should undergo a maintenance. A maintenance mechanism can only lock a region that is not already locked, which prevents different maintenance mechanisms from interfering with each other.
Lock Region Size. A lock region consists of a fixed number of consecutively-addressed DRAM rows. To simplify the design, we set the lock region size to one or multiple subarrays. This is because a maintenance operation uses the local row buffers in a subarray, and thus having a subarray shared by multiple lock regions will cause a conflict when accessing a row in a different lock region that maps to the subarray under maintenance. We design and evaluate SMD assuming a default lock region size of 512-row subarrays.
Modern DRAM chips typically use the density-optimized open-bitline architecture [84, 64], which places sense amplifiers on both top and bottom sides of a subarray and adjacent subarrays share sense amplifiers. With the open-bitline architecture, the MC should not access a row in a subarray adjacent to one under maintenance. To achieve this, the Lock Controller simply sends an ACT_NACK when the MC attempts to activate a row in a subarray adjacent to a locked region. Consequently, when SMD locks a region that spans 512-row subarrays, it prevents the MC from accessing subarrays in a bank with rows ( subarrays). In the folded-bitline architecture [84, 64], adjacent subarrays do not share sense amplifiers with each other. Therefore, SMD prevents access only to the subarrays in a locked region. We evaluate SMD using the density-optimized open-bitline architecture.
Ensuring Timely Maintenance Operations. A maintenance mechanism cannot lock a region with an active row. To lock the region, the maintenance mechanism waits for the MC to precharge the row. DRAM standards specify the maximum allowed time for a row to remain active as 9x tREFI [78, 83, 85]). As such, a maintenance mechanism is delayed at most by 9x tREFI.
4.3 Controlling an SMD Chip
While introducing changes in the memory controller (MC) to handle ACT_NACK, SMD also simplifies MC design and operation as the MC no longer implements control logic to periodically issue DRAM maintenance commands. For example, the MC does not 1) prepare a bank for refresh by precharging it, 2) implement timing parameters relevant to refresh, and 3) issue REF commands. The MC still maintains the bank state information (e.g., whether and which row is open in a bank) and respects the DRAM timing parameters associated with commands for performing access (e.g., ACT, PRE, RD, WR).
@@ 4.3 Controlling an SMD Chip
Handling an ACT_NACK Signal. Fig. 3 depicts a timeline that shows how the MC handles an ACT_NACK signal. Upon receiving an ACT_NACK, the MC waits for ACT Retry Interval (ARI) time to re-issue the same ACT. It keeps re-issuing the ACT command once every ARI until the DRAM chip accepts the ACT. To ensure that a DRAM chip cannot lock a region for prolonged periods, SMD enforces a minimum delay of ARI between the end of a maintenance operation and the start of the next maintenance operation targeting the same region. This allows the MC to serve a request that targets the locked region immediately after the maintenance operation in this region ends. While waiting for ARI, the MC can activate a row from a different lock region or bank, to overlap the ARI latency with a useful operation.
@@
Setting the ACT Retry Interval (ARI). Setting ARI to a very low or high value relative to the expected duration of the maintenance operations can have a negative impact on system performance and energy efficiency. We empirically find that is a favorable configuration for the three maintenance mechanisms that we evaluate in this work (section 5).
@@
Latency. An SMD chip sends an ACT_NACK DRAM command bus cycles after receiving the ACT. should be low so that the MC is quickly notified when an ACT fails, and the MC can attempt to activate a different row in the same bank while waiting for ARI.
@@
We estimate the overall latency based on the latency of an RD in a conventional DRAM chip. An RD has a latency breakdown that resembles the latency breakdown of a . An RD command 1) propagates from the MC to the DRAM chip, 2) accesses data in a portion of the row buffer in the corresponding bank, and 3) sends the data back to the MC. In the DDR4 standard [78], the latency between issuing an RD and the first data beat appearing on the data bus is defined as tCL (typically 22 cycles for DDR4-3200). The latency components 1) and 3) of RD are similar to those of . Thus, the main difference between and tCL arises from the second component. According to our evaluation, the latency of accessing the Lock Region Table (LRT) is (section 7.4). Given the relatively low complexity of the LRT compared to the datapath that is involved during an RD, we believe the overall latency can be designed to be much smaller than tCL. We assume cycles unless stated otherwise. In our evaluations, we find that small latencies (e.g., ) have negligible effect on system performance mainly because the number of rejected ACTs constitute a small portion of all ACTs.
4.4 ACT_NACK Divergence Across DRAM Chips
@@
SMD maintenance operations take place independently in each DRAM chip. Therefore, when a DRAM rank, which can be composed of multiple DRAM chips operating in lock step, receives an ACT command, some of the chips may send an ACT_NACK while others do not. Normally, ACT_NACK divergence would not happen for maintenance mechanisms that perform the exact same operation at the exact same time in all DRAM chips (e.g., the Fixed Rate Refresh mechanism, Section 5.1). However, a mechanism can also operate differently in each DRAM chip, e.g., a variable rate refresh mechanism [17, 15, 38]. As a result of this divergence, the row becomes partially open in chips that do not send ACT_NACK.
As a solution, the MC waits for ARI and re-issues the same ACT command to retry activating the row in the chips that previously sent an ACT_NACK signal. 555SMD retries activating the DRAM row in chips that rejected the previous activate command by sending a new activate command to all DRAM chips in the rank. DRAM chips that already have an active row simply ignore the activate command. Our analysis shows that under the worse-case ACT_NACK divergence (i.e., when each of the 16 chips in a rank perform refresh on a different region at different time), SMD-VR performs 15.4% worse than SMD-VR with no ACT_NACK divergence, which still provides 7.8% average speedup over conventional DDR4 refresh for memory-intensive four-core workloads.
4.5 Ensuring Forward Progress for Memory Requests
The MC could fail to retry an ACT_NACK’d memory request if an active DRAM row (e.g., in a lock region other than the one targeted by the rejected memory request) precludes retrying the ACT_NACK’d memory request every ARI. Such retry failures could occur repeatedly and lead to temporary starvation of a lock region (i.e., no request targeting that lock region is served for a prohibitively long time). Temporary starvation of a lock region is very unlikely to happen because it requires a maintenance mechanism to repeatedly lock the same lock region for maintenance, which is not a property of a carefully designed maintenance mechanism. Instead, a carefully designed maintenance mechanism alternates between different lock regions every as it performs new maintenance operations (e.g., periodic refresh section 5.1).
To prevent temporary starvation of lock regions irrespective of maintenance mechanism design, SMD issues an ACT_NACK’d memory request every ARI. No memory request is ACT_NACK’d by two separate maintenance operations because an SMD chip does not perform any maintenance operations for ARI after the end of a maintenance operation in a lock region (section 4.3). Issuing a rejected memory request strictly every ARI does not prevent the MC from serving memory requests that access other lock regions. Within one ARI () the MC can 1) activate a DRAM row other than the ACT_NACK’d DRAM row, 2) serve seven (based on the standard values of , , timing parameters) READ/WRITE requests to this newly activated row, and 3) precharge the bank such that an activate command to the previously ACT_NACK’d DRAM row can be issued every ARI.
Proof of Forward Progress. We consider a memory request in the memory controller’s request queue to have made forward progress if it is served by the memory controller.
Theorem. If the MC retries every ARI then makes forward progress.
Proof. We devise a contrapositive proof for the above theorem. Suppose does not make forward progress, i.e., is never dequeued from the request queue. must target a locked region because otherwise is dequeued when it is scheduled. The maintenance mechanism can only hold the lock for finite time. When the maintenance mechanism releases the lock, no maintenance mechanism locks the same region for at least ARI. If the MC retries within one ARI after the lock is released, then is not rejected by a maintenance mechanism, i.e., makes forward progress. However, if the MC retries every ARI, then the MC also retries within one ARI after the lock is released. Thus, if does not make forward progress, then the MC does not retry every ARI. We infer from this that the theorem is true.
Memory Request Stall Time. To limit the maximum potential memory request stall time, 1) long time taking maintenance operations can be divided into many smaller short time taking maintenance operations, 2) the maximum time a maintenance operation takes can be specified in the DRAM standard. In our design, the most time taking maintenance operation is ECC scrubbing (Section 5.3), which takes to read a DRAM row. This stall time is comparable to what a memory request in a modern DDR4-based system experiences. In such a system, a request that arrives at the memory controller’s request queues immediately after the memory controller schedules a periodic refresh operation (by issuing a command) has to wait for the refresh operation to complete (e.g., for [78]).
4.6 Impact to Request Scheduling
The key drawback of our SMD implementation is that it makes the synchronous DDRx interface less predictable. However, the practical performance overheads of this drawback are very low, as we demonstrate in Section 7.